OpenStudy (anonymous):

solve the system by using elimination or elimination with multiplication. Write the solution as an ordered pair. (x'y) a-b=-5 a+2b=1

OpenStudy (anonymous):

use same approach as above to get one constant as negative so that it can cancel out a - b = -5 and -1(a + 2b = 1) a-b=-5 -a-2b=-1 add them together--> -3b=-6, div both sides by -3, so b=2 plug b back into the 1st equation, so a-2=-5, a=-3 so (a,b) = (-3,2)

OpenStudy (anonymous):

hope that helps u!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcome!

OpenStudy (anonymous):

a=-3 and b=2